# Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection

### v2

@@ Abstract
Most toxicity detection models treat toxicity as an intrinsic property of text, overlooking the role of context in shaping its impact. Drawing on interdisciplinary research, we reconceptualise toxicity as a socially emergent stress signal. We introduce a new framework for toxicity detection, including a formal definition and metric, and validate our approach on a novel dataset, demonstrating improved contextual sensitivity and adaptability.
Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection
@@ 1 Introduction
When asked to define obscenity, judge Potter Stewart famously said: “I know it when I see it". To this day, we have not come far from this definition.
Natural language processing (NLP) has made remarkable technical progress in detecting “toxic” language. Yet, despite ever-larger models and increasing benchmark scores, a foundational question remains unresolved: what exactly is toxicity?
An influential study by Dixon et al. (2018), defines toxic speech as “rude, disrespectful, or unreasonable language likely to make someone leave a discussion.” However, terms such as “rude” and “disrespectful” are highly subjective, varying considerably across cultures, individuals, and situational contexts. Crucially, such definitions offer no quantitative basis for measurement, leading to labelling inconsistencies that undermine model performance Garg et al. (2023). Despite all its vagueness and flaws, this definition is still widely used by researchers and industry.
Such ambiguity prevents consistent classification of even simple phrases - for instance, “I don’t like you”. Some may interpret it as toxic, others may not. Asking annotators with differing perspectives to label toxicity leads to disagreement, introducing noise that undermines model accuracy Liu et al. (2021); Goyal et al. (2022); Sap et al. (2021).
At the level of academia as a whole, the lack of a standardised definition has led researchers to employ different terms for similar work Vidgen (2019), assigning labels of “toxic", “hateful", “offensive", “abusive", etc. to effectively the same data Fortuna (2020); Madukwe et al. (2020). Considering this, in this paper, we use “toxic speech” as an umbrella term.
Another major issue in toxicity detection research is the field’s overreliance on benchmark-driven evaluation. Improvements in benchmark scores are frequently mistaken for real-world progress, despite the absence of ground truth metrics linking these scores to reductions in social harm such as exclusion, stress, or violence.
Most toxicity detection systems implicitly assume that toxicity is an intrinsic property of language - something that can be detected from surface features, isolated utterances, or predefined keyword lists. This assumption neglects the crucial role of context. Without context, toxicity detection systems are prone to racial, sexual, political, religious, and geographical biases Garg et al. (2023); Sap et al. (2019). These systems often degrade into simple profanity filters Chen (2022), lacking the nuance necessary for real-world use Xenos et al. (2021); Pavlopoulos et al. (2020); Menini et al. (2021) and becoming susceptible to numerous adversarial attacks Berezin et al. (2025, 2024, 2023).
Regarding the issues described above, researchers often argue that improving results requires a bigger and better model. Contrary to that, we suggest that we do not need a bigger model but a better definition.
@@ 1.1 Our Contribution: A New Frame, A Measurable Signal
We propose a paradigm shift: from viewing toxicity as an intrinsic property of language to understanding it as an emergent signal of perceived social stress. Drawing interdisciplinary research, we frame toxicity as a contextual stress reaction - one that becomes evident not solely through the content of a message, but through its social reception.
To operationalise this perspective, we introduce PONOS (Proportion Of Negative Observed Sentiments), a novel metric that quantifies toxicity based on the collective emotional response a message provokes. Rather than determining what a message intrinsically "is", PONOS captures how it is received.
2 Critics of current definitions
Contrary to the longstanding adage - "sticks and stones may break my bones, but words will never hurt me", contemporary research says: words can inflict measurable psychological and physiological harm Struiksma et al. (2022b); Sapolsky (2004).
Toxic communication has been linked to increased anxiety, chronic stress, radicalisation, and - at its extreme - self-harm and suicide van Geel et al. (2014); Kiritchenko et al. (2021); Cervone et al. (2021). At the societal level, hate speech exacerbates discrimination, fosters violence, and erodes the conditions for open discourse Vidgen and Derczynski (2021).
Yet despite the urgency of addressing these harms, the field lacks a coherent operational definition of toxicity. Definitions across academic, legal, and industrial contexts remain fragmented – often vague, inconsistent, and mutually incompatible Mnassri et al. (2024).
In an effort to systematise existing approaches, Herz and Molnár (2012) categorise definitions along four dimensions:
@@ 2 Critics of current definitions
Each category, however, reveals critical shortcomings:
@@
Harm-based definitions rely on an inherently vague notion of harm, which encompass both genuine toxicity and benign but distressing acts (e.g., delivering bad news).
@@
Content-based definitions risk overreach, potentially classifying legitimate political disagreement as hate speech Parekh (2012); Hare and Weinstein (2009); Anderson and Barnes (2023).
@@
Lexical-based approaches overlook toxic messages that use ostensibly neutral language (e.g., “Women, back to the kitchen!”).
@@
Dignity-based models conflate legitimate critique with toxicity, risking the suppression of vital social and political discourse Anderson and Barnes (2023).
@@
Subsequent work by Khurana et al. (2022) identifies five critical dimensions – target group, dominance, perpetrator identity, type of negative reference, and type of consequences – but finds no universal definition applicable across datasets or social contexts. Even expanded notions like Fiser et al.’s “socially unacceptable discourse” Fišer et al. (2017) fail to resolve the underlying contextual instability.
@@
As Brown (2017) aptly observes, “hate speech” is not merely contested but “systematically ambiguous,” carrying different meanings depending on the cultural, political, and historical frames invoked.
3 Recontextualizing Toxicity: An Interdisciplinary Framework for Verbal Harm
To move beyond the limitations of current toxicity detection approaches, we synthesise insights from multiple fields to develop a framework for understanding toxic speech.
3.1 Personal level
The experience of verbal harm is fundamentally affective, rooted in how individuals perceive threats to personal and social stability. Research in psychology and neuroscience agrees that such harm does not stem from the intrinsic properties of language, but from context-dependent appraisals, emotional expectations, and violations of predictive frameworks.
In psychology appraisal theory Lazarus (1991) posits that emotional and stress responses arise when individuals evaluate events as threatening to their goals or social standing. In this view, communication becomes harmful not because of its content alone, but because it is perceived as violating significant personal concerns.
Research on social pain shows that symbolic threats can activate biological stress systems in much the same way as physical injury Eisenberger et al. (2003); Lieberman and Eisenberger (2009).
Neuroscientific findings indicate that the brain processes language in an anticipatory, predictive, and context-sensitive manner. Predictive coding models Friston (2010); Kleckner et al. (2017); Barrett and Simmons (2015) suggest that the brain constantly generates expectations about emotional and social states. When communication violates these expectations, it triggers stress responses.
Physiological indicators of stress, such as heart rate variability (HRV), are influenced by exposure to social threats Shaffer et al. (2014). ERP studies Ding et al. (2016); Liang et al. (2018) demonstrate that negative emotional language disrupts neural integration, while chronic exposure to social stressors leads to pro-inflammatory genetic activity Slavich and Cole (2013).
Together, this research shows that stress responses are closely tied to exposure to harmful or toxic speech, and that the resulting harm is not merely psychological - it is physiological as well.
3.2 Group level
Beyond direct victimization, witnessing harm to others can also trigger stress responses, signalling potential threats within the social environment Struiksma et al. (2022a).
Sociology and social psychology demonstrate that responses to verbal behaviour are governed by group norms, emotional expectations, and mechanisms of norm enforcement Elster (2007). Research by Bicchieri (2005) and Cialdini et al. (1991) shows that social norms are maintained through collective emotional and behavioural reactions to deviance. These violations are not assessed in isolation, but rather interpreted through the emotional dynamics of the group, which shape perceptions of harm. The concepts of emotional collectives van Kleef and Fischer (2015) and emotional norm-shaping processes van Kleef (2024); Sarı et al. (2024) emphasize how shared emotions reinforce or reshape social expectations in real time.
Moral Foundations Theory further underscores that emotional responses to norm violations differ across groups and cultures Graham et al. (2013)
From this perspective, toxicity functions as a social signal - a disruption in the moral and emotional fabric of the group.
3.3 Cultural-Language level
Linguistic and sociolinguistic research demonstrates that meaning is co-constructed in interaction, subject to contextual framing. Harmful speech cannot be reduced to individual words or expressions divorced from their social use.
Pragmatics Clark (1996); Sperber and Wilson (1986/1995) emphasises that meaning is not fully encoded in words but inferred through shared background knowledge and conversational context. An identical phrase may be benign in one setting and deeply offensive in another, depending on tone, power dynamics, and prior discourse.
Politeness theory and impoliteness theory Brown and Levinson (1987); Culpeper (1996) show that perceived rudeness, offence, or aggression are not intrinsic to utterances, but arise from how speech acts align (or misalign) with social expectations. In this view, offence is an emergent property of interaction — not a lexical attribute.
Language ideologies Woolard and Schieffelin (1994); Irvine (2001) highlight that judgements about what counts as “toxic” or “neutral” language are shaped by power structures and dominant cultural narratives.
Historical linguistics and semantics further demonstrates the change of meaning over the change of context Allan and Burridge (2006); Traugott and Dasher (2001). Words such as “nice” (originally meaning “foolish”) or “fag” (a benign term in British English, a slur in American usage) reveal how emotional and moral associations shift across cultural and historical contexts. This semantic fluidity undermines any fixed, universal standard of toxicity.
Multilingualism and language variation add further complexity. Emotional resonance and perceived offensiveness vary across languages due to differing socialisation patterns Dewaele (2004). What is taboo, rude, or inappropriate in one language and culture may carry no such connotation in another.
Together, these insights argue against treating harmful language as a static phenomenon. Toxicity cannot be reliably detected by identifying “bad words” in isolation, but only through an understanding of the forces that shape how language functions in context.
@@ 4 Definition of Toxicity
Building on those insights, we identify two core components of verbal harm: (1) context-dependent violation of social or moral norms, and (2) the induction of stress and emotional dysregulation as a consequence.
For example, consider a speaker who writes a profane, intentionally offensive message in Chinese to a recipient who does not speak the language. Although the content is objectively toxic and the intent is hostile, no harm is experienced if the recipient cannot interpret the message. There is no perceived norm violation, no affective response, and no stress. Under our framework, such a message would be classified as low in toxicity - not because it is morally defensible, but because it fails to generate a measurable signal of harm. The message would be toxic if it reached an informed audience.
Conversely, imagine a speaker makes a flirtatious joke, believing it to be benign or even complimentary. If the recipient experiences it as inappropriate or violating - due to context, power dynamics, or personal boundaries - stress and social disruption will occur. In many professional and legal contexts, this is recognised as harm: offence is taken, not given. Under our model, such an interaction would register as toxic, even in the absence of harmful intent.
These cases illustrate a crucial point: toxicity is defined not by intent or language alone, but by its reception - by the stress it induces and the norms it is perceived to violate.
Based on this synthesis, we offer the following definitions:
@@
Toxicity - the characteristic of causing stress through contradiction of accepted morality and norms of interaction within the context of interaction.
@@
Toxic speech - speech that causes stress through contradiction of accepted morality and norms of communication within the context of interaction.
@@
These definitions deliberately foreground the role of context, perception, and reception. We distinguish between toxicity and toxic speech to account for the broader scope of harmful communicative acts, including visual, multimodal, or behavioural signals that may not be strictly linguistic.
@@
It is important to clarify that while our framework uses stress as a measurable proxy for toxicity, we do not claim that all stress-inducing communication is inherently toxic. For instance, delivering bad news or asserting uncomfortable truths may elicit stress without violating social or moral norms. Our definition of toxicity focuses on the intersection of perceived norm violation and stress response, not on stress in isolation. This distinction is crucial to avoid conflating necessary discomfort with communicative harm.
@@
This reconceptualisation challenges the assumption that toxicity can be located in isolated words or surface features. Instead, it positions toxicity as a socially emergent signal - a disruption to shared expectations that results in emotional or relational stress. Any detection system that hopes to mitigate real-world harm must take this interactional, context-sensitive perspective into account.
@@
We are not denying intent’s relevance, but argue that it’s insufficient alone and serves rather as a part of a context. In our definition we use context as the full set of conditions under which a communicative act is interpreted - including verbal, situational, cognitive, cultural, and other socially relevant dimensions. The more precisely we understand these conditions, the more accurately we can assess the acceptability. A more detailed explanation can be found in Appendix A.
4.1 Towards Measurement: Stress as an Objective Metric
If the goal of toxicity detection is to mitigate harm, then we must define harm in terms that are both scientifically grounded and computationally tractable. Across disciplines, stress consistently emerges as a robust and measurable proxy for communicative harm, particularly in cases where harm is symbolic, ambiguous, or socially mediated.
Unlike subjective judgments of offensiveness or rudeness, stress is a quantifiable physiological state, akin to body temperature or blood pressure McEwen (1998); Thayer et al. (2012). It offers a principled anchor for modelling emotional impact, enabling us to move beyond vague or culturally biased notions of toxicity.
Stress can be inferred through a spectrum of physiological and behavioural indicators, each offering trade-offs between accuracy and scalability:
Direct physiological measures, such as glucocorticoid levels (e.g., cortisol), offer high biological accuracy and are widely used in laboratory studies. However, these require invasive sampling (blood, saliva, urine) and are not viable for real-time or large-scale deployment.

### v3

# From Intrinsic Toxicity to Reception-Based Toxicity: A Contextual Framework for Prediction and Evaluation

@@ Abstract
Most toxicity detection models treat toxicity as an intrinsic property of text, overlooking the role of context in shaping its impact. In this position paper, drawing on insights from psychology, neuroscience, and computational social science, we reconceptualise toxicity as a socially emergent signal of stress. We formalise this perspective in the Contextual Stress Framework (CSF), which defines toxicity as a stress-inducing norm violation within a given context and introduces an additional dimension for toxicity detection. As one possible realisation of CSF, we introduce PONOS (Proportion Of Negative Observed Sentiments), a metric that quantifies toxicity through collective social reception rather than lexical features. We validate this approach on a novel dataset, demonstrating improved contextual sensitivity and adaptability when used alongside existing models.
From Intrinsic Toxicity to Reception-Based Toxicity: A Contextual Framework for Prediction and Evaluation
@@ 1 Introduction
When asked to define obscenity, Judge Potter Stewart famously said: "I know it when I see it". To this day, we have not come far beyond this intuition.
Natural language processing (NLP) has made remarkable technical progress in detecting "toxic" language. Yet, despite ever-larger models and rising benchmark scores, a foundational question remains unresolved: what exactly is toxicity?
An influential work by def_old defines toxic speech as "rude, disrespectful, or unreasonable language likely to make someone leave a discussion". However, terms such as "rude" and "disrespectful" are inherently subjective, varying considerably across cultures, individuals, and situational contexts. Crucially, such definitions offer no quantitative basis for measurement, leading to labelling inconsistencies that undermine model performance survey.
Despite its vagueness and limitations, this definition remains widely used in both research and industry.
This ambiguity prevents consistent classification even for seemingly simple phrases - for instance, "I don’t like you". Some may interpret it as toxic, while others may not. Asking annotators with differing perspectives to label toxicity therefore leads to disagreement, introducing noise that undermines model accuracy authors1; autors2; authors3.
At the level of academia as a whole, the lack of a standardised definition has led researchers to employ different terms for similar phenomena vidgen-etal-2019-challenges, assigning labels such as "toxic", "hateful", "offensive", or "abusive" to effectively the same data its_all_same; madukwe-etal-2020-data. In this paper, we use toxic speech as an umbrella term.
Another major issue in toxicity detection research is the field’s overreliance on benchmark-driven evaluation. Improvements in benchmark scores are frequently treated as evidence of real-world progress, despite the absence of ground-truth metrics linking these scores to reductions in social harm, such as exclusion, stress, or violence.
Most toxicity detection systems implicitly assume that toxicity is an intrinsic property of language - something that can be detected from surface features, isolated utterances, or predefined keyword lists. This assumption neglects the crucial role of context. Without context, toxicity detection systems are prone to racial, sexual, political, religious, and geographical biases survey; bias_nig. These systems often degrade into simple profanity filters holy_shit, lacking the nuance necessary for real-world use xenos-etal-2021-context; context2; context3 and becoming susceptible to numerous adversarial attacks tip; berezin2024; berezin-etal-2023-offence.
@@ 1.1 Our Contribution: A New Frame, A Measurable Signal
To address this, we argue for a shift in emphasis: from viewing toxicity as purely intrinsic to the text towards modelling it as an emergent signal of perceived social stress.
Drawing on interdisciplinary insights, we introduce the Contextual Stress Framework (CSF), which frames toxicity as a stress-inducing norm violation within a given context. CSF adds an additional dimension to existing text-intrinsic approaches by capturing how toxicity is perceived and socially received.
Within CSF, we introduce PONOS (Proportion Of Negative Observed Sentiments) as one possible realisation of the framework: a metric that quantifies toxicity via the collective emotional reception a message provokes. Rather than determining what a message intrinsically is, PONOS captures how it is received in context.
2 Related Work and Limitations of Existing Definitions
Contemporary research demonstrates that words can inflict measurable psychological and physiological harm struiksma2022people; sapolsky2004zebras.
Toxic communication has been linked to increased anxiety, chronic stress, radicalisation, and, at its extreme, self-harm and suicide VanGeel2014; Kiritchenko2021; Cervone2021. At the societal level, hate speech exacerbates discrimination, fosters violence, and erodes the conditions for open public discourse Vidgen2020.
Despite the urgency of addressing these harms, the field lacks a coherent operational definition of toxicity. Definitions across academic, legal, and industrial contexts remain fragmented, often vague, inconsistent, and mutually incompatible mnassri2024survey.
In an effort to systematise existing approaches, Herz2012 categorise definitions of harmful speech along four dimensions:
@@ 2 Related Work and Limitations of Existing Definitions
Each category, however, exhibits critical limitations.
@@
Harm-based definitions rely on an inherently underspecified notion of harm, encompassing both genuinely toxic speech and socially necessary but distressing acts (e.g. delivering bad news).
@@
Content-based definitions risk overreach, potentially classifying legitimate political disagreement or critique as hate speech parekh_2012; hare; sep-hate-speech.
@@
Lexical-based approaches overlook toxic messages expressed through ostensibly neutral language (e.g. "Women, back to the kitchen!").
@@
Dignity-based models risk conflating legitimate criticism with toxicity, thereby suppressing vital social and political discourse sep-hate-speech.
@@
Subsequent work by khurana-etal-2022-hate identifies five critical dimensions: target group, dominance, perpetrator identity, type of negative reference, and type of consequences. Yet, it finds no universal definition applicable across datasets or social contexts. Even broader formulations, such as Fiser et al.’s notion of "socially unacceptable discourse" fiser, fail to resolve the underlying contextual instability.
@@
As Brown2017 observes, harmful speech is not merely contested but systematically ambiguous, acquiring different meanings depending on the cultural, political, and historical frames in which it is situated.
3 An Interdisciplinary Basis for CSF
Across psychology, neuroscience, and sociolinguistics, verbal harm is understood as fundamentally context-dependent: the same utterance can be harmless or harmful depending on who interprets it, under what expectations, and within which normative setting. A common thread is an appraisal process: communicative acts are evaluated relative to local moral and interactional norms, and when an act is perceived as threatening, disrespectful, or boundary-violating, it can trigger affective responses and stress-related dysregulation in recipients and bystanders lazarus1991emotion; eisenberger2003does; lieberman2009neuroscience.
At the individual level, research on stress and emotion emphasises that harmful impact depends on interpretation rather than surface form alone. Predictive and context-sensitive accounts of cognition model perception as expectation-driven; violations of anticipated social or emotional states can elicit stress responses even when intent is ambiguous friston2010free; barrett2015interoceptive; kleckner2017evidence. This link between social threat and physiological regulation is reflected in established correlates such as heart-rate variability shaffer2014healthy.
At the interactional level, harm is jointly constructed and becomes legible through sequences of action and reaction: norms are maintained and renegotiated through observable responses such as condemnation, withdrawal, or escalation. Work on norm enforcement shows that deviance is regulated through collective emotional and behavioural reactions Bicchieri_2005; CIALDINI1991201; vankleef2015emotions.
At the cultural level, communities differ in how they sanction profanity, sarcasm, reclaimed slurs, or in-group teasing; what appears "toxic" out of context may be benign in-group signalling, and vice versa. Pragmatic accounts emphasise that meaning is inferred from shared background and interactional framing rather than encoded in words alone clark1996using; Sperber1995-SPER. Politeness and impoliteness research likewise treat offence as an emergent property of interaction Brown_Levinson_1987; CULPEPER1996349, while language ideology work highlights how judgements of "toxic" versus "neutral" language are shaped by power and dominant cultural narratives woolard1994language; irvine2001style. Finally, meaning and offensiveness shift across time and across languages, undermining any fixed universal standard Traugott_Dasher_2001; Dewaele01062004.
Taken together, these literatures motivate CSF’s core composition: toxicity arises when an utterance is perceived to violate local norms and this perception induces stress or disruption in an audience (additional discussion in Appendix A).
@@ 4 Definition of Toxicity
Building on the interdisciplinary insights outlined above, we model verbal harm as comprising two components: (i) a perceived violation of social or moral norms in a given interactional context; and (ii) an ensuing stress response in those who interpret the act.
Consider a speaker who writes a profane, intentionally offensive message in Chinese to a recipient who does not speak the language. Although the intent is hostile and the string contains profanity, no harm is experienced if the recipient cannot interpret it. There is no perceived norm violation, no affective response, and no stress. Under our framework, such a message registers as low in toxicity in this interaction - not because it is morally defensible, but because it fails to produce a measurable signal of harm in the audience that actually receives it. The same message would become toxic if it reached an informed audience.
Conversely, imagine a speaker makes a flirtatious joke believing it to be benign or even complimentary. If the recipient experiences it as inappropriate or violating - due to context, power dynamics, or personal boundaries, stress and social disruption may occur. In many professional and legal contexts, this is recognised as harm: offence is taken, not given. Under our model, such an interaction can be toxic even in the absence of harmful intent.
These examples motivate a crucial point: toxicity is defined not by intent or language alone, but by its reception - by the stress it induces and the norms it is perceived to violate.
Based on this synthesis, we propose:
@@
Toxicity — the characteristic of causing stress via a perceived contradiction of accepted morality and norms of interaction, within the context in which it is interpreted.
@@
Toxic speech — speech that induces stress via a perceived contradiction of accepted morality and norms of communication, within the context in which it is interpreted.
@@
These definitions deliberately foreground context, perception, and reception. We distinguish between toxicity and toxic speech to account for the broader scope of harmful communicative acts, including visual, multimodal, or behavioural signals that may not be strictly linguistic.
Operational implication (two-axis view). The two components of the definition above yield two complementary measurement routes. Text-only toxicity detectors (e.g. moderation models trained on annotator judgements) primarily operationalise perceived norm violation from the content itself under assumed or averaged contexts. In contrast, reception-based signals operationalise experienced stress as it manifests in interaction (e.g. in replies), capturing context that is not present in the message text alone. Our framework therefore treats intrinsic scores and reception signals as distinct axes rather than interchangeable "toxicity scores": an utterance may be norm-violating without eliciting strong negative reception (high violation, low stress), or elicit negative reception without being flagged by text-only detectors (low violation, high stress).
@@
While our framework treats stress as a measurable proxy for harm, we do not claim that all stress-inducing communication is inherently toxic. Delivering bad news or asserting uncomfortable truths may elicit stress without violating social or moral norms. Toxicity, as defined here, arises from the intersection of perceived norm violation and stress response, not from stress in isolation. This distinction is crucial to avoid conflating necessary discomfort with communicative harm.
@@
This reconceptualisation challenges the assumption that toxicity can be located in isolated words or surface features. Instead, it positions toxicity as a socially emergent signal - a disruption of shared expectations that manifests as emotional or relational stress. Any detection system intended to mitigate real-world harm must therefore adopt an interactional, context-sensitive perspective.
@@
We do not deny the relevance of speaker intent, but argue that intent alone is insufficient and should instead be treated as one component of context. Here, context refers to the full set of conditions under which a communicative act is interpreted, including linguistic, situational, cognitive, cultural, and relational dimensions. A more detailed discussion of contextual factors is provided in Appendix B.
4.1 Towards Measurement: Stress as a Measurable Signal
If toxicity detection aims to mitigate harm, then the relevant notion of harm must be grounded in a signal that is (i) theoretically motivated and (ii) measurable at scale. Across disciplines, stress provides such an anchor: it refers to a family of physiological and behavioural responses that arise when an individual appraises a situation as threatening, boundary-violating, or socially unsafe. This makes stress a principled correlate of communicative harm in cases where harm is symbolic, ambiguous, or socially mediated.
Unlike global labels such as "offensive" or "rude", stress has a well-characterised biological basis and a long measurement tradition in psychology and neuroscience mcewen1998; thayer2012.